OpenStudy (nubeer):

lim x-->0 [(xe^(1/x))/(lnx^sinx)] anyone plz can help me on this question..

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{e^{\frac{1}{x}}}{\ln(x)^{\sin(x)}}\]??

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{xe^{\frac{1}{x}}}{\ln(x)^{\sin(x)}}\]

OpenStudy (anonymous):

tried using l'hopital's rule and it just go way uglier. the wolf says one way it is zero and the other way is infinity http://www.wolframalpha.com/input/?i=limit+as+x+goes+to+0+%28xe%5E%281%2Fx%29%29%2F%28ln%28x%29%5Esin%28x%29%29